OpenStudy (anonymous):

If 4^y=7, then the value of 4^3y+2 is?

OpenStudy (ash2326):

\[4^y=7\] \[4^{3y}=(4^y)^3\] So you'll get \[\large 4^{3y}=7^3\] Can you solve now?

OpenStudy (anonymous):

343 Thank you:)

jimthompson5910 (jim_thompson5910):

343 isn't the final answer, but it gets you there

OpenStudy (anonymous):

343+2=345

jimthompson5910 (jim_thompson5910):

is the +2 in the exponent when you say 4^3y+2 ie is the problem \[\Large 4^{3y+2}\] or is it \[\Large 4^{3y}+2\]

jimthompson5910 (jim_thompson5910):

if it's the second one, then you're correct

OpenStudy (anonymous):

It's the second one 4^(3y)+2

jimthompson5910 (jim_thompson5910):

ok then you got it, just making sure

OpenStudy (anonymous):

Nice, thank you!